Sin theta = 4/5 then find the value of cosec theta /1+cot theta

Grade:11

4 Answers

Arun
5 years ago
Dear student
 
sin theta = 4 / 5
hence
 
cosec theta = 5/4
&
cot theta = ¾
 
now
cosec theta / (1+ cot theta))
 = (5/4) / (1+3/4)
 = (5/4) / (7/4)
 = 5/7
 
Regards
Arun (askIITians forum expert)
Shraddha
5 years ago
Sorry sir I have a doubt regarding this when I solved I got the same answer as yours but in my book 20/21 is the answer given
Aarushi Ahlawat
5 years ago
There are two possible answers for the given questions but none of those is 20/21.
 
One is 5/7 and other one 5.
 
For sin(theta)=4/5
 
cot(theta)= either 3/4 or -3/4 .
 
So possible answers are 5/7 and 5.  
Aarushi Ahlawat
5 years ago
20/21 is one of the answer for sec(theta)/(1+cot(theta) other possible answer for this is -20/3. There may be typo mistake either in question or answer.